### v1

@@ Abstract
Understanding the origin and evolution of supermassive black holes (SMBH) stands as one of the most important challenges in astrophysics and cosmology, with little current theoretical consensus. Improved observational constraints on the cosmological evolution of SMBH demographics are needed. Here we report results of a search via photometric variability for SMBHs appearing as active galactic nuclei (AGN) in the cosmological volume defined by the Hubble Ultra Deep Field (HUDF). This work includes particular focus on a new observation carried out in 2023 with the Hubble Space Telescope (HST) using the WFC3/IR/F140W, which is compared directly to equivalent data taken 11 years earlier in 2012. Two earlier pairs of observations from 2009 to 2012 with WFC3/IR/F105W and WFC3/IR/F160W are also analysed. We identify 443, 149, and 78 AGN candidates as nuclear sources that exhibit photometric variability at a level of 2, 2.5 and 3 $\sigma$ in at least one filter. This sample includes 29, 14, and 9 AGN at redshifts $z>6$, when the Universe was $\lesssim 900$ Myr old. After variability and luminosity function (down to $M_{\rm UV}=-17\>$mag) completeness corrections, we estimate the co-moving number density of SMBHs, $n_{\rm SMBH}(z)$. At $z=6-9$, $n_{\rm SMBH}\gtrsim 10^{-2}\>{\rm cMpc^{-3}}$. At low-$z$ our observations are sensitive to AGN fainter than $M_{\rm UV}=-17\>$mag, and we estimate $n_{\rm SMBH}\gtrsim 6\times 10^{-2}\>{\rm cMpc^{-3}}$. We discuss how these results place strong constraints on a variety of SMBH seeding theories.
@@ 1 Introduction
In Paper I of this series (Hayes et al., 2024) we reported first results from a new observation of the HUDF that was made in 2023 in the F140W filter (GO 17073, PI: Hayes, hereafter HUDF23). The images were taken to an equivalent depth as the HUDF12 observation, and thus permit a detailed study of variability between the two epochs. An analysis of variability between the HUDF09 and HUDF12 epochs was also carried out. The main result of Paper I was the report of three high-$z$, i.e., between $z=6-7$, AGN candidates identified via variability and implication of these results for $n_{\rm SMBH}$ and thus SMBH seeding models.
@@ 2.3 Identifying Variable Sources
Following Paper I, we identify variables in each matched pair of filters using two techniques: comparing the nuclear/central photometry of galaxies at different epochs and detecting residual sources in pair-subtracted images in any given filter. Regions near the image edges are excluded due to the dithering pattern causing excess noise and we focus solely on a central region of the HUDF covering 123″$\times$139″.
@@ 2.3.1 Photometric Variables
[…]
@@
To calibrate our estimates of the uncertainties, we compare the observed standard deviations of the $\Delta m$ distributions in magnitude bins of 0.50.50.5\>0.5mag width over the range from 24 to 31 mag with the estimated mean photometric uncertainties in each bin from Eq. (1). We then compute the total average scale factor that globally adjusts the uncertainty reported in Eq. (1) to match the observed $\Delta m$ standard deviation. For the F105W, F140W, and F160W filters the resulting calibration scale factors are 0.80, 0.84, and 0.79, respectively. We refer to the calibrated uncertainties as $\delta_{\Delta m,{\rm cal}}$.
Next we carry out a linear fit to the logarithm of the calibrated $1\sigma$ uncertainty estimates in each 0.5 mag bin. It is by a comparison of this estimated $1\sigma$ uncertainty as a function of magnitude that a given source’s significance of variability is ultimately assessed in a two-epoch observation in a given filter.
We consider three thresholds of significance for variability: $2\sigma$; $2.5\sigma$; and $3\sigma$. For a source to be classified as a candidate variable, it must have a $\sigma_{\Delta m}\equiv\Delta m/\delta_{\Delta m,{\rm cal}}\geq 2$ in at least one filter.
@@ 2.3.2 Difference Imaging Variables
As a complement to variability identified via aperture photometry, we also generate pair-subtracted images for each filter. From these images we may identify variable sources that were not detected via aperture photometry. We are also able to assess the morphology of photometric variables. In particular, single AGN are expected to appear as compact, unresolved variable sources. For this process, each image is first smoothed with a small Gaussian kernel of 0.5 pixels. Then images taken in the same filters were directly subtracted, resulting in six “difference images”. We then applied Source Extractor to these difference images and effective gains were re-calculated using the difference of the squared exposure times to account for the increased sky noise in the subtracted images, which mimics a shallower integration time. Each detected source is manually inspected, and spurious artefacts are removed. Difference image selected sources, as we describe below, are relatively few in number, but are considered to have high significance and for the purposes of counting statistics are included in the $3\sigma$ significance photometric sample.
@@ 2.3.3 Redshift determination
We cross-correlate each analyzed source with positions from known redshift catalogs, including spectroscopic data from VLT/MUSE (Bacon et al., 2023), the JADES GTO program for spectroscopic redshifts using NIRSpec (Bunker et al., 2023a), photometric redshifts from NIRCam + HST (Rieke et al., 2023), and the Ultraviolet UDF photometric catalog (Rafelski et al., 2015). Spectroscopic estimates, when available, are given preference over photometric ones.
@@ 3 Results
Figure 1: (a) Top panel: Photometric variability of HUDF sources in the F140W filter from 2012 to 2023, $\Delta m_{\rm F140W}$, versus mean magnitude, $\bar{m}_{\rm F140W}$ (gray points). Black circles show the intrinsic $1\sigma$ scatter of the data measured in bins of 0.5 mag width. Yellow circles depict the calibrated (by a factor of 0.84) photometric $1\sigma$ uncertainties in each bin (see text). The inner pair of yellow lines are fits to these calibrated uncertainties. The next sets of yellow lines show $2\sigma$ and $3\sigma$ uncertainties, scaled from the $1\sigma$ fit. Red squares highlight sources that are estimated to be $\geq 3\sigma$ variables. (b): Bottom panel: As (a), but now showing the variability significance metric, $\sigma_{\Delta m,{\rm F140W}}$, versus $\bar{m}_{\rm F140W}$. A linear scale is used for $|\sigma_{\Delta m,{\rm F140W}}|\leq 3$ and a logarithmic scale elsewhere. The cross-matched known AGNs in the HUDF (Lyu et al., 2022) are marked by open squares. Sources identified by difference imaging are shown in open circles.
@@ 3.1 Detected Variable Sources
Figure 1b shows the dimensionless variability metric $\sigma_{\Delta m,{\rm F140W}}$ of each source versus mean magnitude. A linear scale is used for the inner range from $-3$ to $+3$ and then a logarithmic scaling at higher absolute values. The limits of $\sigma_{\Delta m,{\rm F140W}}=2,2.5,3$ are highlighted, showing the different levels of significance that we consider. From these data we identify 219, 74 and 41 sources that are $>2\sigma$, $>2.5\sigma$, and $>3\sigma$ variables in the F140W observations (see Table 3).
@@
The equivalent results of Figure 1 for the F105W and F160W observations are shown in Appendix A. From these data, the equivalent numbers of variable sources from the F105W (2009-2012, i.e., 3-year time baseline) observation are 247, 112 and 54 sources at $2\sigma$, $2.5\sigma$, and $3\sigma$ variables, respectively. The F160W (2009-2012, i.e., 3-year time baseline) observation yields 218, 83 and 39 sources at $2\sigma$, $2.5\sigma$, and $3\sigma$ variables, respectively. We see that the F105W, F140W and F160W observations have similar efficiencies at detecting variable sources.
The difference imaging method identifies a further 18 variable sources, which were not detected via the photometric method. In the following we consider these sources as part of the $\geq 3\sigma$ sub-sample.
Combining the above results for all the filters, in the end we obtain a total sample of 596, 233, $106+19=124$ sources that show variability in at least one filter at the $2\sigma$, $2.5\sigma$, and $3\sigma$ levels, respectively. Of these, 553, 211, and 92 sources have been cross-matched with measured redshift catalogues, respectively, with the redshift distribution discussed in the next sub-section. The photometric variability data for each source with a known redshift is presented in the electronic version of Table 2, with the displayed version here listing only the highest redshift ($z>6$) sources.
@@ 3.2 Redshift Distribution
Figure 2 presents a scatter plot of $\sigma_{\Delta m}$ versus $z$ for the F105W, F140W, and F160W datasets. We see that the redshift distribution of the sources in the sample extends out to $z\sim 9.5$, with the highest redshift $3\sigma$ variability detections being at $z\sim 8$ (two sources) and $2.5\sigma$ detections being at $z=8.3$ and 8.7. The latter is the highest $z$ variable in our sample. Table 2 lists all the 25 variable sources that we identify with $z>6$ (note the online electronic table contains the full sample at all redshifts).
Figure 3a shows the redshift histogram of our sample of detected variable sources. Different colors depict the different samples according to the significance level ($>2$, 2.5 and $3\sigma_{m}$) as detected in at least one filter. We see that most candidates reside at $z\lesssim 4$ for all cases of significance threshold. Figure 3b shows the fraction of galaxies that are detected to have variable nuclei. This fraction appears relatively constant with redshift and is seen to be elevated compared to the false positive levels, with this enhancement relatively greater at higher values of $\sigma$. However, the absolute number of variable sources are often dominated by those found at $2\sigma$, even after subtracting off the number of expected false positives (see next sub-section).
For a visual representation of the high redshift variable sources in the field, Figure 4 shows their positions superimposed on the HUDF image. Different symbols indicate the different ranges of significance level in units of $\sigma$ as detailed in the legend, color coded according to the estimated redshift. Zoom-in images of the highest redshift ($z>8$) variable candidates are presented in small inset panels. We note that the brightest of these, source id 1829, has a nearby, lower redshift (photometric $z=2.25$) present in its inset.
Figure 4: Angular distribution of variable candidates in the HUDF at $z\geq 6$. The estimated redshift is shown by the color of the symbol with reference to the color-bar. Different symbols refer to different significance levels (see legend). We also highlight sources at $z>8$ with zoom-in thumbnails.
@@
3026 53.15032 -27.77523 $29.74\pm 0.40$ $30.15\pm 0.35$ $29.94\pm 0.36$ $30.01\pm 0.56$ $29.69\pm 0.26$ $30.23\pm 0.44$1.74 -0.31 -2.71 3.24 $8.74^{+0.62}_{-0.58}$
785 53.14414 -27.79393 $30.18\pm 0.55$ $30.38\pm 0.41$ $30.16\pm 0.47$ $29.93\pm 0.51$ $30.17\pm 0.35$ $30.79\pm 0.49$0.67 0.81 -2.16 2.40 $8.16^{+0.03}_{-0.10}$
2617 53.14410 -27.77863 $29.55\pm 0.32$ $29.93\pm 0.28$ $30.23\pm 0.45$ $29.85\pm 0.55$ $30.35\pm 0.56$ $29.65\pm 0.26$1.91 1.40 3.39 4.13 $8.09^{+6.93}_{-0.14}$
1829 53.16415 -27.78452 $29.57\pm 0.50$ $29.10\pm 0.13$ $29.09\pm 0.15$ $28.97\pm 0.20$ $28.95\pm 0.13$ $29.02\pm 0.15$ 3.84 1.12 -0.81 3.47 $8.02^{0.23}_{-0.11}$
2165 53.16824 -27.78230 $30.00\pm 0.50$ $30.44\pm 0.43$ $30.53\pm 0.59$ $30.51\pm 0.86$ $30.74\pm 0.66$ $30.05\pm 0.39$1.50 -0.02 2.32 2.77 $7.53^{+0.46}_{-0.29}$
3326 53.16745 -27.77203 $30.08\pm 0.49$ $30.81\pm 0.58$ $30.86\pm 0.68$ $31.18\pm 1.65$ $30.88\pm 0.72$ $30.16\pm 0.42$2.30 -0.58 2.16 3.21 $7.274$b
3351 53.16382 -27.77174 $30.37\pm 0.62$ $30.49\pm 0.39$ $30.19\pm 0.39$ $31.19\pm 1.15$ $30.20\pm 0.39$ $30.20\pm 0.41$0.38 -2.24 -0.00 2.27 $7.27^{+0.11}_{-0.75}$
2567 53.18626 -27.77897 $29.31\pm 0.31$ $29.26\pm 0.13$ $28.79\pm 0.10$ $28.81\pm 0.13$ $28.70\pm 0.09$ $28.85\pm 0.10$ 0.57 -0.12 -2.12 2.20 $7.27^{+0.11}_{-0.75}$
1297 53.16481 -27.7882 $29.18\pm 0.21$ $29.19\pm 0.12$ $28.71\pm 0.09$ $28.52\pm 0.12$ $28.70\pm 0.08$ $28.60\pm 0.09$ 0.14 2.49 1.40 2.86 $7.227$c
2533 53.14653 -27.77936 $29.70\pm 0.36$ $30.19\pm 0.34$ $29.98\pm 0.37$ $29.61\pm 0.37$ $29.72\pm 0.25$ $29.94\pm 0.34$2.01 1.66 -1.25 2.96 $7.2^{+1.0}_{-5.2}$
3270 $5$27.77267 $29.57\pm 0.34$ $30.02\pm 0.29$ $29.60\pm 0.26$ $30.14\pm 0.70$ $29.67\pm 0.28$ $29.76\pm 0.30$2.20 -2.36 -0.60 3.29 $7.17^{+0.09}_{-0.07}$
2247 53.17686 -27.78148 $30.89\pm 1.05$ $30.51\pm 0.49$ $30.38\pm 0.60$ $31.54\pm 1.67$ $30.64\pm 0.66$ $30.53\pm 0.64$ 0.96 -2.06 0.22 2.28 $7.02^{+0.12}_{-0.03}$
3790 53.15560 -27.76475 $30.74\pm 0.81$ $29.99\pm 0.28$ $29.87\pm 0.32$ $29.77\pm 0.44$ $30.02\pm 0.31$ $29.72\pm 0.27$ 2.51 0.43 1.61 3.02 $7.00^{+2.97}_{-5.83}$
2875 53.18089 -27.77668 $30.34\pm 0.67$ $29.84\pm 0.31$ $29.63\pm 0.25$ $29.37\pm 0.36$ $29.72\pm 0.30$ $29.75\pm 0.30$ 2.26 1.51 -0.18 2.73 $6.91^{+0.03}_{-0.10}$
1906 53.14554 -27.78385 $28.17\pm 0.08$ $28.22\pm 0.05$ $28.14\pm 0.05$ $28.20\pm 0.09$ $28.06\pm 0.05$ $28.25\pm 0.07$0.60 -0.87 -4.07 4.20 $6.874$c
1511a 53.16193 -27.78699 $28.48\pm 0.16$ $28.22\pm 0.06$ $28.05\pm 0.07$ $27.77\pm 0.08$ $27.83\pm 0.05$ $27.99\pm 0.06$ 5.04 6.32 -4.25 9.13 $6.74^{+0.04}_{-0.07}$
2709 53.17761 -27.77821 $30.98\pm 1.26$ $30.74\pm 0.62$ $31.17\pm 1.19$ $30.30\pm 0.78$ $30.53\pm 0.60$ $31.72\pm 1.83$ 0.54 1.75 -2.30 2.95 $6.72^{+0.27}_{-0.25}$
1058 53.15794 -27.79093 $30.03\pm 0.65$ $30.73\pm 0.40$ $29.96\pm 0.32$ $30.28\pm 0.71$ $30..6\pm 0.29$ $30.14\pm 0.39$2.33 -1.15 -0.37 2.62 $6.56^{+0.001}_{-6.37}$
101159d 53.16052 -27.78593 $28.96\pm 0.37$ $\gtrsim 30.1$ $28.96\pm 0.37$ $\gtrsim 30.1$ $\gtrsim 29.9$ $\gtrsim 29.9$. $6.54^{+2.45}_{-2.59}$
93 53.16556 -27.80509 $30.61\pm 1.08$ $30.39\pm 0.63$ $30.75\pm 0.95$ $30.68\pm 1.54$ $30.70\pm 1.07$ $30.08\pm 0.64$ 0.67 0.09 2.07 2.17 $6.53^{+0.44}_{-4.83}$
305 53.15643 -27.80044 $30.02\pm 0.49$ $29.96\pm 0.30$ $30.51\pm 0.65$ $29.87\pm 0.48$ $30.23\pm 0.45$ $30.17\pm 0.42$ 0.48 2.09 0.28 2.16 $6.26^{+0.19}_{-0.02}$
3391 53.16101 -27.77124 $29.92\pm 0.41$ $29.76\pm 0.21$ $29.78\pm 0.31$ $30.50\pm 0.67$ $29.89\pm 0.28$ $30.00\pm 0.32$ 0.97 -2.48 -0.53 2.71 $6.21^{+0.31}_{-0.05}$
261 53.15983 -27.80118 $30.36\pm 0.67$ $30.39\pm 0.45$ $30.44\pm 0.56$ $30.75\pm 1.33$ $32.13\pm 2.60$ $30.66\pm 0.63$0.08 -0.80 2.25 2.39 $6.09^{+1.65}_{-2.28}$
2736 53.18742 -27.77797 $29.44\pm 0.47$ $30.53\pm 0.65$ $31.04\pm 1.25$ $31.72\pm 5.86$ $30.93\pm 0.87$ $30.37\pm 0.59$4.82 0.00 1.50 5.05 $6.08^{+0.03}_{-0.03}$
2680 53.15550 -27.77841 $29.97\pm 0.46$ $30.33\pm 0.39$ $31.43\pm 1.45$ $30.33\pm 0.74$ $30.09\pm 0.38$ $30.99\pm 0.89$1.28 1.97 -2.92 3.75 $6.01^{+0.05}_{-0.06}$
@@
Table 2: General properties of the $z>6$ variable sources at $>2\sigma$ significance (bold text) combining the three different filters. aThis source corresponds to source 1052123 in Paper I. bSpectroscopic redshift from NIRspec (Bunker et al., 2023b). cSpectroscopic redshift from FRESCO (Oesch et al., 2023). 1Comparing epochs taken in 2008-9 vs 2012. 2Comparing epochs taken in 2012 vs 2023. dThis source is taken from Hayes et al. (2024) and was only detected via image subtraction method as described in the text.

### v2

@@ Abstract
Understanding the origin and evolution of supermassive black holes (SMBH) stands as one of the most important challenges in astrophysics and cosmology, with little current theoretical consensus. Improved observational constraints on the cosmological evolution of SMBH demographics are needed. Here we report results of a search via photometric variability for SMBHs appearing as active galactic nuclei (AGN) in the cosmological volume defined by the Hubble Ultra Deep Field (HUDF). This work includes particular focus on a new observation carried out in 2023 with the Hubble Space Telescope (HST) using the WFC3/IR/F140W, which is compared directly to equivalent data taken 11 years earlier in 2012. Two earlier pairs of observations from 2009 to 2012 with WFC3/IR/F105W and WFC3/IR/F160W are also analysed. We identify 521, 188, and 109 AGN candidates as nuclear sources that exhibit photometric variability at a level of 2, 2.5 and 3 $\sigma$ in at least one filter. This sample includes 13, 3, and 2 AGN candidates at redshifts $z>6$, when the Universe was $\lesssim 900$ Myr old. After variability and luminosity function (down to $M_{\rm UV}=-17\>$mag) completeness corrections, we estimate the co-moving number density of SMBHs, $n_{\rm SMBH}(z)$. At $z\gtrsim 6$, $n_{\rm SMBH}\gtrsim 6\times 10^{-3}\>{\rm cMpc^{-3}}$. At low-$z$ our observations are sensitive to AGN fainter than $M_{\rm UV}=-17\>$mag, and we estimate $n_{\rm SMBH}\gtrsim 10^{-2}\>{\rm cMpc^{-3}}$. We discuss how these results place strong constraints on a variety of SMBH seeding theories.
@@ 1 Introduction
In Paper I of this series (Hayes et al., 2024) we reported first results from a new observation of the HUDF that was made in 2023 in the F140W filter (GO 17073, PI: Hayes, hereafter HUDF23). The images were taken to an equivalent depth as the HUDF12 observation, and thus permit a detailed study of variability between the two epochs. An analysis of variability between the HUDF09 and HUDF12 epochs was also carried out. The main result of Paper I was the report of three high-$z$, i.e., between z=67z=6-7, AGN candidates identified via variability and implication of these results for $n_{\rm SMBH}$ and thus SMBH seeding models.
@@ 2.3 Identifying Variable Sources
[…]
@@ 2.3.1 Photometric Variables
$\delta_{\Delta m}=\Big{(}\frac{2.5}{{\rm ln}\>10}\Big{)}\sqrt{\Big{(}\frac{\delta_{F_{\mathrm{1}}}}{F_{1}}\Big{)}^{2}+\Big{(}\frac{\delta_{F_{\mathrm{2}}}}{F_{2}}\Big{)}^{2}},$ (1)
@@
To calibrate our estimates of the uncertainties, we compare the observed standard deviations of the $\Delta m$ distributions in magnitude bins of 0.50.5\>mag width over the range from 24 to 31 mag with the estimated mean photometric uncertainties in each bin from Eq. (1).
We then compute the total average scale factor that globally adjusts the uncertainty reported in Eq. (1) to match the observed $\Delta m$ standard deviation. For the F105W, F140W, and F160W filters the resulting calibration scale factors are 0.80, 0.84, and 0.79, respectively. We refer to the calibrated uncertainties as $\delta_{\Delta m,{\rm cal}}$.
It is by taking the ratio of the $\Delta m$ with the uncertainty calculated in Eq. 1 that a given source’s significance of variability is ultimately assessed in a two-epoch observation in a given filter. We verify that this empirical approach follows the trend of the calibrated uncertainties by carrying out a linear fit to the logarithm of the calibrated $1\sigma$ uncertainty estimates in each 0.5 mag bin (see Fig. 1). We consider three thresholds of significance for variability: $2\sigma$; $2.5\sigma$; and $3\sigma$. For a source to be classified as a candidate variable, it must have a $\sigma_{\Delta m}\equiv\Delta m/\delta_{\Delta m,{\rm cal}}\geq 2$ in at least one filter.
@@ 2.3.2 Difference Imaging Variables
As a complement to variability identified via aperture photometry, we also generate pair-subtracted images for each filter. From these images we may identify additional variable sources, including sources offset from galactic nuclei, that were not detected as variables via aperture photometry. Also, for extended sources we are able to assess the morphology of photometric variables. In particular, single AGN are expected to appear as compact, unresolved variable sources. Then images taken in the same filters were directly subtracted, resulting in six “difference images”. We then applied Source Extractor to these difference images and effective gains were re-calculated using the difference of the squared exposure times multiplied by the CCD gain to account for the increased sky noise in the subtracted images, which mimics a shallower integration time. Difference image selected sources, as we describe below, are relatively few in number, but are considered to have high significance and for the purposes of counting statistics are included in the $3\sigma$ significance photometric sample.
Finally, to prevent possible image artefacts and other forms of interlopers contaminating our sample, we visually inspect every source identified as variable in both methods. Possible interlopers include moving objects, noisy pixels, saturated sources, and telescope diffraction spikes. SNe which occur between the two epochs (either between 2009 and 2012 or between 2012 and 2023) are removed from the sample since there will be no detection in the first epoch and a bright, obvious detection in the second. For this visual inspection, along with pair-subtracted images, we check the single epoch frames in all the three filters. Furthermore, we rule out the possibility of low redshift interlopers or artefacts by examining cutouts in the optical and in the infrared by leveraging available observations with HST and JWST, respectively (see Fig. 5 for the used criteria). Sources identified as image artefacts or interlopers are excluded from the variable source sample.
@@ 2.3.3 Redshift determination
We cross-correlate each analyzed source with positions from known redshift catalogs, including spectroscopic data from VLT/MUSE (Bacon et al., 2023), the JADES GTO program for spectroscopic redshifts using NIRSpec (Bunker et al., 2023), photometric redshifts from NIRCam + HST (Rieke et al., 2023), and the Ultraviolet UDF photometric catalog (Rafelski et al., 2015). Spectroscopic estimates, when available, are given preference over photometric ones. We note that the average photometric redshift uncertainty is of the order of $\lesssim 0.2$, which justifies our choice of unit width bins when considering the overall redshift distribution (see nest Section).
@@ 3 Results
Figure 1: (a) Top panel: Photometric variability of HUDF sources in the F140W filter from 2012 to 2023, $\Delta m_{\rm F140W}$, versus mean magnitude, $\bar{m}_{\rm F140W}$ (gray points). Black circles show the intrinsic $1\sigma$ scatter of the data measured in bins of 0.5 mag width. Yellow circles depict the calibrated (by a factor of 0.84) photometric $1\sigma$ uncertainties in each bin (see text). The inner pair of yellow lines are fits to these calibrated uncertainties. The next sets of yellow lines show $2\sigma$ and $3\sigma$ uncertainties, scaled from the $1\sigma$ fit. Red squares highlight sources that are estimated to be $\geq 3\sigma$ variables. (b): Bottom panel: As (a), but now showing the variability significance metric, $\sigma_{\Delta m,{\rm F140W}}$, versus $\bar{m}_{\rm F140W}$. A linear scale is used for $|\sigma_{\Delta m,{\rm F140W}}|\leq 3$ and a logarithmic scale elsewhere. The cross-matched known AGNs in the HUDF (Lyu et al., 2022) are marked by open squares. Variable sources identified only by difference imaging are shown in open circles.
@@ 3.1 Detected Variable Sources
Figure 1b shows the dimensionless variability metric $\sigma_{\Delta m,{\rm F140W}}$ of each source versus mean magnitude. A linear scale is used for the inner range from 3-3 to +3+3 and then a logarithmic scaling at higher absolute values. The limits of $\sigma_{\Delta m,{\rm F140W}}=2,2.5,3$ are highlighted, showing the different levels of significance that we consider. From these data we identify 191, 67 and 39 sources that are $2\sigma$, $2.5\sigma$, and $3\sigma$ variables in the F140W observations (see Table 3).
@@
The equivalent results of Figure 1 for the F105W and F160W observations are shown in Appendix A. From these data, the equivalent numbers of variable sources from the F105W (2009-2012, i.e., 3-year time baseline) observation are 222, 92 and 43 sources at $2\sigma$, $2.5\sigma$, and $3\sigma$ variables, respectively. The F160W (2009-2012, i.e., 3-year time baseline) observation yields 185, 68 and 39 sources at $2\sigma$, $2.5\sigma$, and $3\sigma$ variables, respectively. We see that the F105W, F140W and F160W observations have similar efficiencies at detecting variable sources.
The difference imaging method identifies a further 6 variable sources, which were not detected as variable via the photometric method. In the following we consider these sources as part of the $\geq 3\sigma$ sub-sample. Additionally, 3 more sources previously identified at $2\sigma$ and 1 source identified at $2.5\sigma$ are promoted to $3\sigma$ for a total of 10 sources added at this significance level.
Combining the above results for all the filters, in the end we obtain a total sample of 521, 188, $99+10=109$ sources that show variability in at least one filter at the $2\sigma$, $2.5\sigma$, and $3\sigma$ levels, respectively. Of these, 482, 173, and 95 sources have been cross-matched with measured redshift catalogues, respectively, with the redshift distribution discussed in the next sub-section. The photometric variability data for each source with a known redshift is presented in the electronic version of Table 2, with the displayed version here listing only the highest redshift ($z>6$) sources.
@@ 3.2 Redshift Distribution
Figure 2 presents a scatter plot of $\sigma_{\Delta m}$ versus $z$ for the F105W, F140W, and F160W datasets. We see that the redshift distribution of the sources in the sample extends out to $z\sim 9.5$, with the highest redshift (at $2.5\sigma$) variability detections being at $z\sim 8$ (three sources). Among them, the highest redshift variable in our sample results at $z=7.53^{+0.46}_{-0.29}$. Table 2 lists all the 13 variable sources that we identify with $z>6$ (note the online electronic table contains the full sample at all redshifts).
Figure 3a shows the redshift histogram of our sample of detected variable sources. Different colors depict the different samples according to the significance level (>2>2, 2.5 and $3\sigma_{m}$) as detected in at least one filter. We see that most candidates reside at $z\lesssim 4$ for all cases of significance threshold. Figure 3b shows the fraction of galaxies that are detected to have variable nuclei. This fraction appears relatively constant with redshift and is seen to be elevated compared to the false positive levels, with this enhancement relatively greater at higher values of $\sigma$. However, the absolute number of variable sources are often dominated by those found at $2\sigma$, even after subtracting off the number of expected false positives (see next sub-section).
For a visual representation of the high redshift variable sources in the field, Figures 4 and 5 show their positions superimposed on the HUDF image and cutout frames in different bands used for the visual inspection, respectively. Different symbols indicate the different ranges of significance level in units of $\sigma$ as detailed in the legend, color coded according to the estimated redshift. Zoom-in images of the highest redshift ($z>7$) variable candidates are presented in small inset panels in Figure 4.
Figure 4: Angular distribution of variable candidates in the HUDF at $z\geq 6$. The estimated redshift is shown by the color of the symbol with reference to the color-bar. Different symbols refer to different significance levels (see legend). We also highlight sources at $z>7$ with zoom-in thumbnails.
Figure 5: Cutout images of the sources listed in Table 2. In particular for each source, from left to right, we have F105W difference image, F140W difference image, median ACS image (F606W + F775W + F814W for $z>6.9$ or F435W + F606W + F775W for $z<6.9$) for identifying drop-outs/interlopers, the median F105W image (median stack of both epochs), median F140W image and JADES F277W image (deepest JADES filter image).
@@
2165 53.16823 -27.78223 $30.00\pm 0.19$ $30.44\pm 0.17$ $30.53\pm 0.24$ $30.51\pm 0.35$ $30.74\pm 0.28$ $30.05\pm 0.16$ -1.58 -0.02 2.03 2.57 $7.53^{+0.46}_{-0.29}$
3326 53.16744 -27.77195 $30.08\pm 0.21$ $30.81\pm 0.24$ $30.86\pm 0.32$ $31.18\pm 0.64$ $30.88\pm 0.32$ $30.16\pm 0.17$ -2.31 -0.49 1.85 3.0 7.2747.274b
2567 53.18626 -27.77897 $29.31\pm 0.13$ $29.26\pm 0.06$ $28.79\pm 0.05$ $28.81\pm 0.07$ $28.70\pm 0.05$ $28.85\pm 0.06$ 0.49 -0.13 -2.20 2.26 $7.27^{+0.11}_{-0.75}$
1297 53.16481 -27.7882 $29.18\pm 0.09$ $29.19\pm 0.05$ $28.71\pm 0.05$ $28.52\pm 0.06$ $28.70\pm 0.04$ $28.60\pm 0.04$ 0.16 2.76 1.52 3.16 7.2277.227c
3270 53.16553 -27.77259 $29.57\pm 0.13$ $30.02\pm 0.12$ $29.60\pm 0.10$ $30.14\pm 0.24$ $29.67\pm 0.11$ $29.76\pm 0.12$ -2.34 -2.06 -0.6 3.18 $7.17^{+0.09}_{-0.07}$
1511a 53.16192 -27.78699 $28.48\pm 0.06$ $28.22\pm 0.02$ $28.05\pm 0.02$ $27.77\pm 0.03$ $27.83\pm 0.02$ $27.99\pm 0.02$ 4.58 7.41 -4.8 9.95 $6.74^{+0.04}_{-0.07}$
2080 53.17732 -27.78238 $28.21\pm 0.04$ $28.14\pm 0.02$ $28.04\pm 0.03$ $28.07\pm 0.04$ $28.06\pm 0.03$ $28.01\pm 0.03$ 2.15 -0.51 1.36 2.59 $6.73^{+0.05}_{-0.03}$
1058 53.15794 -27.79093 $30.03\pm 0.19$ $30.73\pm 0.22$ $29.96\pm 0.14$ $30.28\pm 0.27$ $30..6\pm 0.15$ $30.14\pm 0.18$ -2.40 -1.07 -0.36 2.65 $6.56^{+0.001}_{-6.37}$
101159b 53.16052 -27.78593 $28.96\pm 0.37$ $\gtrsim 30.1$ $28.96\pm 0.37$ $\gtrsim 30.1$ $\gtrsim 29.9$ $\gtrsim 29.9$ …. $6.54^{+2.45}_{-2.59}$
305 53.1564 -27.80044 $30.02\pm 0.19$ $29.96\pm 0.11$ $30.51\pm 0.24$ $29.87\pm 0.19$ $30.23\pm 0.18$ $30.17\pm 0.18$ 0.48 2.08 0.27 2.15 $6.26^{+0.19}_{-0.02}$
3391 53.16101 -27.77124 $29.92\pm 0.18$ $29.76\pm 0.09$ $29.78\pm 0.12$ $30.50\pm 0.33$ $29.89\pm 0.13$ $30.00\pm 0.15$ 0.95 -2.01 -0.52 2.29 $6.21^{+0.31}_{-0.05}$
2736 53.18742 -27.77797 $29.44\pm 0.15$ $30.53\pm 0.25$ $31.04\pm 0.50$ $31.72\pm 1.16$ $30.93\pm 0.46$ $30.37\pm 0.29$ -3.67 0.0 0.97 3.79 $6.08^{+0.03}_{-0.03}$
2680 53.15549 -27.77824 $29.97\pm 0.18$ $30.33\pm 0.15$ $31.43\pm 0.53$ $30.33\pm 0.28$ $30.09\pm 0.15$ $30.99\pm 0.37$ -1.36 1.78 -2.31 3.22 $6.01^{+0.05}_{-0.06}$
@@
Table 2: General properties of the $z>6$ variable sources at $2\sigma$ significance (bold text) combining the three different filters. aThis source corresponds to source 1052123 in Paper I. bThis source is taken from Hayes et al. (2024) and was only detected via image subtraction method as described in the text. 1Comparing epochs taken in 2008-9 vs 2012. 2Comparing epochs taken in 2012 vs 2023.